What persent of 240 is 60?




Answer by snwbrdfrk(10)

what you need to do is make that word problem an equation. in this question it will be
? x 240 = 60
you need the question mark to be by itself even though it doesnt have to be a question mark make it what you want. so you need to divide both sides by 240 to get 240 to the other side
? x 240/240 = 60/240
? = 60/240
? = .25
to check your work just type .25(25%)x 240 and you should get 60